OpenStudy (anonymous):

Subtracting the rational expressions x + 1/ 3x +1 minus x/ x- 4 and find the restrictions that resist with (or in) the resulting rational expression

OpenStudy (anonymous):

@ranga @phi @hartnn

OpenStudy (phi):

Please use the equation editor (see button below the input area)

OpenStudy (anonymous):

\[\frac{ 2x }{ 3x+1} - \frac{ x }{ x-4}\]

OpenStudy (phi):

the common denominator will be \[ (3x+1)(x-4) \] what should you multiply the first fraction by to get this common denominator?

OpenStudy (anonymous):

You would have to multiply it by 1, right?

OpenStudy (phi):

btw, your equation does not match what you posted (x+1) in the top of the first fraction versus 2x ??

OpenStudy (anonymous):

I mean't to do 2x, I must of accidentantly hit the one, I'm sorry.

OpenStudy (phi):

before going on, which problem are we doing ?

OpenStudy (anonymous):

The one I wrote with the equation editor. Disregard the x+1 I was typing fast and meant 2x

OpenStudy (anonymous):

@phi

OpenStudy (phi):

ok, so multiply the first fraction by 1, but in the form \[ \frac{(x-4)}{(x-4)} \] and the second fraction by \[ \frac{(3x+1)}{(3x+1)} \]

OpenStudy (phi):

btw, I would *not* multiply out the denominator (x-4)(3x+1) we want it in factored form, because we want to find any x's that make the bottom zero (these will be the x's that are not allowed)

OpenStudy (anonymous):

Okay, and for (X-4) would it be 1(x-4) with the multiplication (sorry it's not in equation editior for some reason it was messing up for me)

OpenStudy (phi):

like this \[ \frac{ 2x }{ (3x+1)}\cdot \frac{(x-4)}{(x-4)} - \frac{ x }{ (x-4)}\cdot \frac{(3x+1)}{(3x+1)} \]

OpenStudy (phi):

distribute the 2x in 2x*(x-4) distribute the -x in -x*(3x+1)

OpenStudy (anonymous):

2x-4 -3x +1 ?

OpenStudy (phi):

For the first fraction 2x * (x-4) distribute the 2x means 2x times each term inside the parens you write 2x*x and 2x * -4 then simplify

OpenStudy (anonymous):

So it would be 2x and -8x?

OpenStudy (phi):

almost, except 2*x * x is 2*x*x (you get x*x in there) people use the short-hand x^2 for x*x (i.e. \(x^2= x\cdot x\))

OpenStudy (anonymous):

So could we use 2x^2

OpenStudy (phi):

yes, the top of the first fraction becomes 2x^2 -8x

OpenStudy (anonymous):

\[\frac{ 2^{x2}-8x }{ (3x+1)(x-4) }\]

OpenStudy (phi):

now the second fraction distribute the -x in -x*(3x+1) -3x^2 -x

OpenStudy (phi):

I assume you mean \[ \frac{ 2x^2-8x }{ (3x+1)(x-4) } \]

OpenStudy (anonymous):

Yes, still getting used to equation editior, i'm sorry!

OpenStudy (phi):

yes, the editor is more tricky than the algebra.

OpenStudy (anonymous):

So the numerator for the second equation is -3x^2- x and now we gotta solve, right?

OpenStudy (phi):

you don't solve fractions. but you can add the tops when the two fractions have the same denominator (as we have here)

OpenStudy (phi):

in other words we now have \[ \frac{ 2x^2-8x -3x^2 -x }{(3x+1)(x-4)} \]

OpenStudy (phi):

using the same rule that allows \[ \frac{1}{4} + \frac{2}{4} = \frac{1+2}{4} \]

OpenStudy (anonymous):

Hmm, okay. So now is it kinda like combining the like terms? so it's 5x^2 - 8x

OpenStudy (phi):

combine like terms 2 x^2 take away 3 x^2 leaves how many x^2 ?

OpenStudy (phi):

-8 x's take away another x gives how many x's ?

OpenStudy (anonymous):

We have -1x^2 and -9x's?

OpenStudy (phi):

yes.

OpenStudy (anonymous):

So the new numerator now is -1x^2 - (-9x)?

OpenStudy (phi):

so that is the top if this is multiple choice, you might have to tweak the answer to match the choices. ** So the new numerator now is -1x^2 - (-9x)? ** we already took care of the minus sign when we distributed -x (instead of just x) so the the top is just \[ -x^2 -9x \]

OpenStudy (anonymous):

It's not multiple choice, it's part of an assigment. THe numerator would be correct.

OpenStudy (phi):

now for the *excluded values* we are not allowed to divide by zero the bottom has (3x+1)(x-4) if either factor is zero, the whole bottom becomes zero (and we do not want that) what x makes x-4 = 0 ? what x makes 3x+1 = 0 ?

OpenStudy (anonymous):

-4 for the first one and -3 for the second?

OpenStudy (phi):

ok on the first. that means x=4 is excluded 3*-3 + 1 is -9+1 = -8 which is not zero. use algebra for the second problem 3x+ 1 = 0 first add -1 to both sides 3x + 1 -1 = 0-1 and simplify 3x = -1 now divide both sides by 3

OpenStudy (phi):

*for the first one, I am sure you mean x=4 makes x -4 = 0

OpenStudy (anonymous):

So this means the "restrictions" are x=4 and x=-3x?

OpenStudy (phi):

how did you get x= -3x ? (which is only true if x is zero) 3x+ 1 = 0 first add -1 to both sides 3x + 1 -1 = 0-1 and simplify 3x = -1 now divide both sides by 3

OpenStudy (anonymous):

Well because negative 1 divided by 3 is a decimal.. so I assumed you would do 3x divided by -1 unless it's a decimal for the restriction

OpenStudy (phi):

3x = -1 now divide both sides by 3 -1 divided by 3 is written -⅓ on the left side 3x/3 is just x (because 3/3 is 1 and 1*x is x) we ge x= -⅓ as your other restricted value.

OpenStudy (anonymous):

Oh i'm sorry, thanks so much though for your help, I really appreciate it